How many grams of each reactant is needed to produce 50 grams of Fe2(SO4)3


1
Expert's answer
2021-02-01T03:47:20-0500

50 grams of Fe2(SO4)3 = "\dfrac{\textsf{mass}}{\textsf{molar mass}} = \dfrac{50}{400} = 0.125 \textsf{ moles}"


"Fe_2(SO_4)_3 = 2Fe + 3S + 3O_4"


This means that one mole of Fe2(SO4)3 has two moles of Iron atom, three moles of Sulphur atom, and 12 moles of Oxygen atom.


Therefore, 0.125 moles of Fe2(SO4)3 has 0.25 moles of Iron atom, 0.375 moles of Sulphur atom, and 1.5 moles of Oxygen atom.


mass of iron atoms in the compound = 0.25 × molar mass = 0.25 × 55.85 = 13.96g

mass of sulfur atoms in the compound = 0.375 × molar mass = 0.375 × 32 = 12g

mass of oxygen atoms in the compound = 1.5 × molar mass = 1.5 × 16 = 24g